OpenStudy (anonymous):

This is a logical question:- If the distance between the Vertex and the focus of a parabola tends to infinity then the parabola assumes the shape of ????

OpenStudy (anonymous):

a line?

OpenStudy (anonymous):

what line

OpenStudy (anonymous):

\[\color{red}{\text{BBUUUUURRRRNNN!!!}}\]

OpenStudy (anonymous):

I think dpalnc is right

OpenStudy (anonymous):

yes a line at least mention that it is straight

OpenStudy (anonymous):

umm, if the vertex is at the origin and and this is an x^2 type parabola, the line y=0 or the x-axis...?

OpenStudy (anonymous):

or a line parallel to the directrix...

OpenStudy (anonymous):

keep it coming Raj... I like these thinking questions...

OpenStudy (anonymous):

yea..dplanc is right. if y= ax^2 then the parabola will tend to x=0 and if it is of the type x=by^2 then it will tend to be y=0.

OpenStudy (anonymous):

I think you mean the otherwise akshat

OpenStudy (anonymous):

oh yes...sorry! :P

OpenStudy (anonymous):

Keep posting such questions man! :D

OpenStudy (anonymous):

will try